# Scaling Robust Optimization for Multi-Agent Robotic Systems: A Distributed Perspective

### v1

@@ III-C2 Distributed Approach
${\bm{\lambda}}_{i}^{\bar{u},l+1}={\bm{\lambda}}_{i}^{\bar{u},l}+\rho_{\bar{u}}(\tilde{{\bm{\mu}}}^{\bar{u},l+1}_{i}-\tilde{{\bm{\nu}}}^{\bar{u},l+1}_{i})$ (64)
${\bm{\lambda}}_{i}^{d,l+1}={\bm{\lambda}}_{i}^{d,l}+\rho_{d}(\tilde{{\bm{\mu}}}^{d,l+1}_{i}-\tilde{{\bm{\nu}}}^{d,l+1}_{i})$
After the global variables update, each agent $i$ would receive the updated global variables ${\bm{\nu}}^{\bar{u},l+1}_{j},{\bm{\nu}}^{d,l+1}_{j}$ from all its neighbor agents $j\in\pazocal{N}_{i}$. Using these received variables, the agent $i$ updates the dual variables ${\bm{\lambda}}_{i}^{\bar{u},l+1},{\bm{\lambda}}_{i}^{d,l+1}$ as per the update step (64). It should be noted that in each update step, all the agents do the update in parallel. Further, the agents need not exchange other system information (such as dynamics parameters or uncertainty set parameters).
@@ IV Simulation Results
In this section, we provide simulation results of the proposed method applied to a multi-agent system with underlying stochastic and deterministic uncertainties. In Section IV-A, the effectiveness of the proposed robust constraints is analyzed under different scenarios. In Section IV-B, the scalability of the proposed distributed framework is illustrated. The agents have 2D double integrator dynamics with the exact dynamic matrices and the uncertainty information provided in Section IV of the SM. In all experiments, we set the time horizon to $T=20$. All experiments are also illustrated through the supplementary video [^1] https://youtu.be/yrDwPwg4WFI.
@@ IV-A1 With deterministic disturbances
(c) Robust case, $k=0$
(d) Robust case, $k=10$
(e) Robust case, $k=15$
(f) Robust case, $k=20$
@@
Figure 2: Twenty agents scenario with circle formation task and obstacle: (a) with robust constraints, (b) without robust constraints, (c)-(f) Snapshots of mean trajectories of agents in robust case.
@@
Here, we compare the results of applying the proposed distributed framework for two cases: robust and non-robust. With non-robust, we refer to the case of finding the optimal control trajectory by only considering standard optimization constraints instead of explicitly accounting for the deterministic uncertainties through robust optimization constraints (robust case). Fig. 1 discloses a two-agent scenario with terminal mean and obstacle avoidance constraints. The plots demonstrate $100$ trajectories of the two agents for different realizations of the disturbances. As shown in Fig. 1a, in the robust case, each agent $i$ successfully avoids the obstacle, and its disturbance-free state mean reaches the target terminal mean ${\bm{\mu}}_{target}^{i}$ with uncertainty in its terminal state mean lying inside the set target bounds (represented by black boxes). On the contrary, for the non-robust case in Fig. 1b, the agents might collide with the obstacle in the center, and the uncertainty in the terminal state mean of each agent might lie outside the target bounds. In Fig. 2, we present a scenario of 20 agents in a circle formation, where each agent must reach the opposite side while avoiding collisions with other agents and an obstacle in the center. Fig. 2a and 2b show $100$ realizations for the robust and non-robust case. Clearly, the robust approach is able to successfully steer the agents to their targets while avoiding collisions, in contrast with the non-robust case where the agents might collide with the obstacle. In Fig. 2c - 2f, we show snapshots of the robust trajectories of the agents for different time instants, demonstrating that the agents do not violate any of the constraints during the tasks. Further, it is worth noting that in the robust case, the agents not only avoid collision but also the uncertainty in the state mean of an agent is reduced when it approaches an obstacle or another agent.
@@ IV-A2 With both deterministic and stochastic disturbances
Further, it is interesting to note the following two observations. From Fig. 3a and 3d, it should be observed that uncertainty in the state trajectory is reduced to be smaller than the uncertainty in the mean in the deterministic case by adding covariance constraints. From Fig. 4a and 4b, it should be observed that the actual trajectories of the agents in the deterministic case avoid the obstacle without enforcing the robust chance constraints. This is because the feedback control gain ${\bf K}$ is used to control the deviations in the state due to the deterministic and stochastic uncertainties. Thus, controlling the deviation in the state due to one type of uncertainty also controls the deviation due to another. This might be useful in cases where one of the uncertainties (deterministic or stochastic) is significantly higher than the other. In such a case, we need not explicitly apply constraints to control the effect of the least pronounced uncertainty.
@@ IV-B Scalability of Proposed Distributed Framework
In our experiments, we noticed that the centralized method with robust terminal mean and initial approach obstacle avoidance constraints alone would fail for a time horizon longer than 15 and a number of agents greater than 3. Fig. 5 shows the result of the proposed distributed framework applied to a large-scale scenario with 100 agents, which verifies the scalability of the proposed approach. The constraints considered are robust terminal mean and inter-agent constraints. Further, Fig. 6 reports the computational times of the proposed framework for an increasing number of agents for scenarios with robust terminal mean and inter-agent constraints. All simulations were carried out in Matlab2022b [17] environment using CVX [15] as the modeling software and MOSEK as the solver [2] on a system with an Intel Corei9-13900HX.
@@ V Conclusion
In this paper, we merge principles from the areas of robust optimization, distribution steering and distributed optimization towards obtaining a scalable safe multi-agent control method under both stochastic and deterministic uncertainties. In particular, we present a distributed robust optimization method that exploits computationally tractable approximations of robust constraints and the separable nature of ADMM in order to significantly reduce the computational demands that such approaches would typically have when applied to multi-agent systems. The results verify that the proposed framework is able to effectively address both types of uncertainty and scale for multi-robot systems with up to a hundred agents.
In future work, we wish to extend the applicability of the proposed method and explore additional robust optimization formulations that would be relevant for robotics problems. In particular, since the current method considers linear time-varying dynamics, it would be possible to extend for nonlinear systems through sequential linearization schemes [32, 34]. Additionally, we intend to further explore suitable robust optimization formulations for multi-agent robotics and investigate connections with shooting methods [18, 41]. Finally, incorporating the current framework into hierarchical distributed optimization schemes [36] would be a promising direction for further increasing the scalability of the method.
Figure 5: Large-scale scenario with 100 agents: Mean trajectories of the agents with robust terminal mean and inter-agent collision avoidance constraints.
Figure 6: Scalability analysis: Computational times of the proposed distributed framework for an increasing number of agents.
@@ Acknowledgments
Supplementary Material
In the following, we provide detailed derivations, proofs and experiment data that have been omitted in the main paper.
VI System Dynamics Compact Form
The system dynamics (1) can be written in a more compact form as follows
where
${\bf G}_{0}^{i}=\begin{bmatrix}{\bf I};&{\bf\Phi}^{i}(1,0);&{\bf\Phi}^{i}(2,0);&\dots;&{\bf\Phi}^{i}(T,0)\end{bmatrix},$ (66b)
${\bf G}_{u}^{i}=\begin{bmatrix}{\bf 0}&{\bf 0}&\dots&{\bf 0}\\ B_{0}^{i}&{\bf 0}&\dots&{\bf 0}\\ {\bf\Phi}^{i}(2,1)B_{0}^{i}&B_{1}^{i}&\dots&{\bf 0}\\ \vdots&\vdots&\vdots&\vdots\\ {\bf\Phi}^{i}(T,1)B_{0}^{i}&{\bf\Phi}^{i}(T,2)B_{1}^{i}&\dots&B_{T-1}^{i}\end{bmatrix},$ (66h)
${\bf G}_{{\bm{w}}}^{i}=\begin{bmatrix}{\bf I}&{\bf 0}&\dots&{\bf 0}\\ {\bf\Phi}^{i}(1,0)&D_{0}^{i}&\dots&{\bf 0}\\ {\bf\Phi}^{i}(2,0)&{\bf\Phi}^{i}(2,1)D_{0}^{i}&\dots&{\bf 0}\\ \vdots&\vdots&\vdots&\vdots\\ {\bf\Phi}^{i}(T,0)&{\bf\Phi}^{i}(T,1)D_{0}^{i}&\dots&D_{T-1}^{i}\end{bmatrix},$ (66n)
${\bf G}_{\zeta}^{i}=\begin{bmatrix}{\bf I}&{\bf 0}&\dots&{\bf 0}\\ {\bf\Phi}^{i}(1,0)&C_{0}^{i}&\dots&{\bf 0}\\ {\bf\Phi}^{i}(2,0)&{\bf\Phi}^{i}(2,1)C_{0}^{i}&\dots&{\bf 0}\\ \vdots&\vdots&\vdots&\vdots\\ {\bf\Phi}^{i}(T,0)&{\bf\Phi}^{i}(T,1)C_{0}^{i}&\dots&C_{T-1}^{i}\end{bmatrix},$ (66t)
with
${\bf\Phi}^{i}(k_{1},k_{2})=A_{k_{1}-1}^{i}A_{k_{1}-2}^{i}\dots A_{k_{2}}^{i}\;\text{for}\;k_{1}>k_{2}.$ (67)
VII Robust Linear Constraints
VII-A Expectation of State
@@ VII-C Proposition 1 Proof
VIII Robust Non-convex Constraints and Distributed Architecture
VIII-A Initial Approach for robust nonconvex constraint
Let us rewrite the robust nonconvex norm-of-mean constraint (33) as follows
@@ VIII-A Initial Approach for robust nonconvex constraint
$\begin{bmatrix}\bm{\Gamma}_{i}^{\mathrm{T}}{\bf F}_{1,k}^{i}({\bf K}^{i})\bm{\Gamma}_{i}+\alpha{\bf S}_{i}&\bm{\Gamma}_{i}^{\mathrm{T}}{\bf F}_{2,k}^{i}({\bf K}^{i},\bar{u}^{i})^{\mathrm{T}}\\ {\bf F}_{2,k}^{i}({\bf K}^{i},\bar{u}^{i})\bm{\Gamma}_{i}&{\bf F}_{3,k}^{i}(\bar{u}^{i})-c_{i}^{2}-\alpha\tau^{i}\end{bmatrix}\succeq 0,$ (92c)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\alpha\geq 0.$ (92d)
Nevertheless, the above constraint would still be a nonconvex constraint, as ${\bf F}_{1,k}^{i}({\bf K}^{i})$, ${\bf F}_{2,k}^{i}({\bf K}^{i},\bar{u}^{i})$, and ${\bf F}_{3,k}^{i}(\bar{u}^{i})$ are nonlinear in ${\bf K}^{i}$, $\bar{u}^{i}$. To address this issue, we first linearize the constraint (89) with respect to ${\bf K}^{i},\bar{u}^{i}$ around the nominal ${\bf K}^{i,l},\bar{u}^{i,l}$, to obtain the following affine robust constraint
@@
VIII-B Proposition 2 Proof
Let us initially rewrite the constraint (40) as follows

### v2

@@ III-C2 Distributed Approach
${\bm{\lambda}}^{i,l+1}_{\bar{u}}={\bm{\lambda}}^{i,l}_{\bar{u}}+\rho_{\bar{u}}(\tilde{{\bm{\mu}}}_{\bar{u}}^{i,l+1}-\tilde{{\bm{\nu}}}_{\bar{u}}^{i,l+1})$ (55)
${\bm{\lambda}}^{i,l+1}_{d}={\bm{\lambda}}^{i,l}_{d}+\rho_{d}(\tilde{{\bm{\mu}}}_{d}^{i,l+1}-\tilde{{\bm{\nu}}}_{d}^{i,l+1})$
After the global variables update, each agent $i$ would receive the updated global variables ${\bm{\nu}}_{\bar{u}}^{j,l+1},{\bm{\nu}}_{d}^{j,l+1}$ from all its neighbor agents $j\in\pazocal{N}_{i}$. Using these received variables, the agent $i$ updates the dual variables ${\bm{\lambda}}^{i,l+1}_{\bar{u}},{\bm{\lambda}}^{i,l+1}_{d}$ as per the update step (55). It should be noted that in each update step, all the agents do the update in parallel.
III-D Computational Complexity Analysis
In this section, we theoretically verify the advantages of the proposed approach in terms of computational complexity. For the following analysis, we consider a multi-agent system of $N$ agents, $n_{\text{obs}}$ obstacles, and with each agent having $n_{\text{inter}}$ neighbors. For simplicity, we consider the case of only having deterministic uncertainty related constraints, but the provided analysis can be readily extended for the case where the constraints (42), (43) and (46) are also added. All results are derived assuming the problems are solved with the interior point method which is considered the state-of-the-art for SOCPs and SDPs [5, 24].
III-D1 Complexity reduction due to constraint approximation
First, we demonstrate the complexity reduction that occurs from the constraint reformulation presented in Section III-A2 compared to the SDP reformulation in III-A2. For that, we provide a comparison of the complexity involved in solving each local sub-problem with the proposed distributed method versus a distributed SDP framework. The latter refers to a distributed framework that involves the initial approach SDP reformulation III-A2 for collision avoidance constraints.
Proposition 4.
The computational complexity for solving each local sub-problem with the distributed SDP framework is
$O\Big{(}L_{\text{IP}}T^{5}(n_{\text{inter}}+n_{\text{obs}})^{3}n_{u_{i}}^{2}n_{x_{i}}^{2}n_{d_{i}}^{2}\gamma_{h}^{2}\Big{)}$ (56)
where $L_{\text{IP}}$ is the number of iterations of the interior point method. With the proposed reformulation, this complexity is reduced to
$O\Big{(}L_{\text{IP}}T^{4}(n_{\text{inter}}+n_{\text{obs}})^{3}n_{u_{i}}^{2}n_{x_{i}}^{2}n_{d_{i}}\gamma_{h}^{2}\Big{)}.$ (57)
Proof.
Provided in Appendix-J. ∎
III-D2 Complexity reduction due to decentralization
Subsequently, we illustrate the computational benefits of the proposed distributed algorithm compared to an equivalent centralized approach. As the local subproblems are the most computationally expensive updates in the distributed ADMM algorithm, its complexity can be expressed as follows.
Corollary 1.
The computational complexity of the proposed distributed algorithm is given by
$O\Big{(}L_{\text{ADMM}}L_{\text{IP}}T^{4}(n_{\text{inter}}+n_{\text{obs}})^{3}n_{u_{i}}^{2}n_{x_{i}}^{2}n_{d_{i}}\gamma_{h}^{2}\Big{)},$ (58)
where $L_{\text{ADMM}}$ is the number of ADMM iterations.
For comparison purposes, we also provide the complexity of an equivalent centralized approach.
Proposition 5.
A centralized approach for solving the multi-agent problem 2 with the proposed reformulated constraints would yield a computational complexity of
$O\Big{(}L_{\text{Lin}}L_{\text{IP}}N^{3}T^{4}(n_{\text{inter}}+n_{\text{obs}})^{3}n_{u_{i}}^{2}n_{x_{i}}^{2}n_{d_{i}}\gamma_{h}^{2}\Big{)}.$ (59)
where $L_{\text{Lin}}$ is the number of outer loops where the non-convex constraints are linearized based on the previous solution.
Proof.
Omitted due to similarity with Proposition 4. ∎
@@ IV Simulation Results
In this section, we provide simulation results of the proposed method applied to a multi-agent system with underlying stochastic and deterministic uncertainties. In Section IV-A, the effectiveness of the proposed robust constraints is analyzed under different scenarios. In Section IV-B, the scalability of the proposed distributed framework is illustrated. The agents have 2D double integrator dynamics with the exact dynamic matrices, and the uncertainty information is provided in Appendix-K. In all experiments, we set the time horizon to $T=20$, except if stated otherwise. For every agent, the neighbor agents are considered to be a set of specified number of agents closest to the agent at the initial state. All experiments are also illustrated through the supplementary video  [^1] https://youtu.be/Q5xghaqt4SQ.
@@ IV-A1 With deterministic disturbances
No. of Obstacles
1 2 3
Initial Reformulation (28) 11.58 22.01 -
Proposed Reformulation 0.73 1.37 1.84
TABLE I: Computational time comparison - Initial Reformulation vs proposed reformulation
Here, we demonstrate the effectiveness of the proposed framework by providing a comparison with the non-robust case, and with the initial reformulation case.
(a) Robust mean trajectories
(b) $k=5$
(c) $k=10$
(d) $k=15$
(e) $k=20$
(f) Distance plot
@@
Figure 2: Twenty agents scenario with circle formation task and 21 obstacles: (a) Robust mean trajectories, (b)-(e) Snapshots of the robust mean trajectories of agents, (f) Distance plot showing the minimum distance of each agent from its nearest neighbor at each time step.
@@
First, we compare the results of applying the proposed distributed framework for two cases: robust and non-robust. With non-robust, we refer to the case of finding the optimal control trajectory by only considering standard optimization constraints instead of explicitly accounting for the deterministic uncertainties through robust optimization constraints (robust case). Fig. 1 discloses a two-agent scenario with terminal mean and obstacle avoidance constraints. The plots demonstrate $100$ trajectories of the two agents for different realizations of the disturbances. As shown in Fig. 1a, in the robust case, each agent $i$ successfully avoids the obstacle, and its disturbance-free state mean reaches the target terminal mean ${\bm{\mu}}_{target}^{i}$ with uncertainty in its terminal state mean lying inside the set target bounds (represented by red boxes). On the contrary, for the non-robust case in Fig. 1b, the agents might collide with the obstacle in the center, and the uncertainty in the terminal state mean of each agent might lie outside the target bounds.
(a) Robust Mean Trajectories
(b) $k=10$
(c) $k=15$
(d) $k=20$
Figure 3: Four agents scenario-1 with non-convex obstacles: (a) Robust mean trajectories, (b)-(d) Snapshots of the robust mean trajectories of agents.
(a) Robust Mean Trajectories
(b) $k=10$
(c) $k=15$
(d) $k=20$
Figure 4: Four agents scenario-2 with non-convex obstacles: (a) Robust Mean Trajectories, (b)-(d) Snapshots of the robust mean trajectories of agents.
Next, we analyze the computational efficiency of the proposed robust reformulated constraints over the initial reformulation (28) by comparing the computational time of solving a single-agent problem using both approaches. The problem involves solving for the agent to satisfy the terminal mean constraints and the obstacle constraints. Table I discloses the optimizer time (in seconds) to solve this problem using each reformulation for an increasing number of obstacles. The initial reformulation fails to provide a solution when there are three or more obstacles, and the time taken with the initial reformulation is significantly longer than the proposed formulation. Moreover, it can be observed that, compared with the initial reformulation, the time taken with the proposed reformulation does not increase significantly as the number of obstacles increases.
We now demonstrate the effectiveness of the proposed framework in complex environments. In each case, a robust mean trajectories plot is included to show that the agents reach the target terminal mean bounds without collision with the obstacles. Fig. 2 shows a scenario with twenty agents and 21 obstacles, where each agent in a circular formation is tasked to reach the opposite side without colliding with obstacles and other agents. We consider a time horizon T = 25. In Fig. 2c - 2e, we show snapshots of the robust trajectories of the agents for different time steps, demonstrating that the agents do not violate any of the constraints during the tasks. Additionally, we provide a distance plot in Fig. 2f showing the minimum distance of each agent from its nearest neighbor ($d_{min}^{i,j}$) at each time step, and the red dashed line representing the collision distance threshold (at 0.25). The nearest neighbor of an agent at any time step is estimated based on the distance between the realizations of the agent and that of its neighbors. It can be observed that the minimum distance decreases first as the agents reach the center and then increases as the agents move away from the center. All the minimum distances lie above the collision threshold, proving no inter-agent collisions occur. Fig. 3 and Fig. 4 demonstrate the performance of the proposed framework in the presence of non-convex obstacles. In Fig. 3 and 4, the non-convex structures are formed using 14 and 16 circular obstacles, respectively. From the snapshots in Fig. 3b - 3d and 4b - 4d, it can be observed that the agents were able to steer through the obstacles without any collision. Further, it is worth noting that in the robust case, the agents not only avoid a collision but also the uncertainty in the state mean of an agent is reduced when it approaches an obstacle or another agent.
IV-A2 With both deterministic and stochastic disturbances
@@
(a) Robust Mean Trajectories
(b) $k=10$
(c) Distance Plot
Figure 7: Fifty agent scenario with six obstacles: (a) Robust mean trajectories of agents in the deterministic case, (b) Snapshot of robust mean trajectories of agents. (c) Distance plot showing the minimum distances of the agent from its nearest neighbor agent at each time step.
@@
Further, it is interesting to note the following two observations. From Fig. 5a and 5d, it should be observed that uncertainty in the state trajectory is reduced to be smaller than the uncertainty in the mean in the deterministic case by adding covariance constraints. From Fig. 6a and 6b, it should be observed that the actual trajectories of the agents in the deterministic case avoid the obstacle without enforcing the robust chance constraints. This is because the feedback control gain ${\bf K}$ is used to control the deviations in the state due to the deterministic and stochastic uncertainties. Thus, controlling the deviation in the state due to one type of uncertainty also controls the deviation due to another. This might be useful in cases where one of the uncertainties (deterministic or stochastic) is significantly higher than the other. In such a case, we need not explicitly apply constraints to control the effect of the least pronounced uncertainty.
(a) Robust mean trajectories
(b) $k=12$
(c) Distance Plot
Figure 8: 164-agent scenario: (a) Mean trajectories of agents in the deterministic case, (b) Snapshot of mean trajectories of agents, (c) Distance plot showing the minimum distances of the agent from its nearest neighbor agent at each time step.
@@ IV-B Scalability of Proposed Distributed Framework
As mentioned, the existing robust trajectory optimization approaches are computationally expensive, even for a single-agent case. In this subsection, we show that the proposed distributed framework enables large-scale robust trajectory optimization. In the following experiments, we consider the deterministic case.
Fig. 7 shows a scenario with 50 agents and six obstacles. Each agent has four neighbors. Fig. 7a and 7b show that all the agents reached the terminal target mean bounds and did not collide with the obstacles. Fig. 7c discloses a distance plot of the minimum distance of each agent from its neighbors, with the red dashed line showing the collision distance threshold (at 0.5). It can be observed that the minimum distances are well above the collision distance threshold, proving that there are no inter-agent collisions. Fig. 8 shows a scenario with 164 agents dispersed in four sets, each of a diamond formation. The task is for the agents to reach the target mean bounds by staying in the diamond formation and avoiding collisions with the neighbor agents. It can be observed that the agents reach the terminal target mean bounds. Further, Fig. 8b - 8c confirm that there are no inter-agent collisions. It is interesting to note from Fig. 7c and 8c that the minimum distances in the final time-step are not the same for all the agents, even if the final target positions are equally distanced. This is due to the dissimilarity in the shape of the state mean realizations of different agents (refer to Section III.D of the supplementary material).
Figure 9: Computational time graph w.r.t the number of obstacles.
Figure 10: Computational time graph w.r.t the number of agents.
Now, we analyse the computational efficiency of the proposed method. First, we provide the computational time graph with respect to the number of obstacles in Fig. 9. There are two significant observations. First, the complexity bounds (57) indicate a cubic dependence of the complexity on the number of obstacles. However, from the computational plot, we can infer that, in practice, the complexity is significantly less than the worst-case bound. Second, the increase in the computational time from five obstacles case to the ten obstacles case is less. When fewer obstacles exist, the significant computational burden would be due to the SOCP constraints (35) on ${\bm{\mu}}_{d}^{i}$ bounds.
Next, we compare the proposed distributed framework with the centralized framework using the proposed reformulated constraints. As we have derived in the earlier section, the complexity of the centralized framework (59) is proportional to the cube of the number of agents, while the complexity of the distributed framework does not directly depend on the number of agents. This is justified by the computational graph in Fig. 10, where we can observe that the centralized framework does not scale favorably.
In all the above cases, the computational time for the distributed framework is estimated as follows. In each ADMM iteration, the maximum optimizer time among all the agents is considered. These per-iteration maximum time values are added over all the ADMM iterations that run until convergence. Further details about the scenarios considered for the creation of the plots can be found in Appendix-N. All simulations were carried out in Matlab2022b [18] environment using CVX [16] as the modeling software and MOSEK as the solver [2] on a system with an Intel Corei9-13900K.
@@ V Conclusion
This paper presents a novel approach for multi-agent control under stochastic and deterministic uncertainty, integrating robust optimization, distribution steering, and distributed optimization principles. We developed computationally tractable approximations of robust constraints that significantly reduce the complexity compared to existing methods without compromising safety. Subsequently, we proposed a distributed framework leveraging the proposed constraint reformulations and the separable nature of ADMM, to solve the computationally challenging problem of multi-agent robust trajectory optimization. We also established rigorous computational complexity bounds for the proposed distributed framework, highlighting the framework’s increased computational efficiency compared to centralized approaches. The simulation results demonstrate the scalability of the proposed framework, successfully handling up to 164 agents, and empirically validate the established complexity bounds. Additionally, we present extensive simulations across diverse scenarios, including non-convex obstacle environments, validating the robustness and computational efficiency of the proposed constraint reformulations.
In future work, we wish to explore new distributed robust optimization algorithms that further enhance the safety capabilities, computational efficiency and applicability of such appraoches in multi-agent robotics problems. In particular, we plan to extend the current method for addressing nonlinear dynamics through considering sequential linearization schemes [35, 39] or combinations with shooting techniques [19, 45]. Additionally, we intend to explore distributed learning-to-optimize approaches for accelerating convergence [40], as well as data-driven techniques for learning the uncertainty sets of the agents [47]. Finally, incorporating the current framework into hierarchical distributed schemes [38] would be a promising direction for further increasing scalability.
@@ Acknowledgments
Appendix A Robust Linear Constraints
A-A Expectation of State
@@ A-C Proposition 1 Proof
Appendix B Robust Non-convex Constraints and Distributed Architecture
B-A Initial Approach for robust nonconvex constraint
Let us rewrite the robust nonconvex norm-of-mean constraint ( 26 ) as follows
@@ B-A Initial Approach for robust nonconvex constraint
$\begin{bmatrix}\bm{\Gamma}_{i}^{\mathrm{T}}{\bf F}_{1,k}^{i}({\bf K}^{i})\bm{\Gamma}_{i}+\alpha{\bf S}_{i}&\bm{\Gamma}_{i}^{\mathrm{T}}{\bf F}_{2,k}^{i}({\bf K}^{i},\bar{{\bm{u}}}^{i})^{\mathrm{T}}\\ {\bf F}_{2,k}^{i}({\bf K}^{i},\bar{{\bm{u}}}^{i})\bm{\Gamma}_{i}&{\bf F}_{3,k}^{i}(\bar{{\bm{u}}}^{i})-c_{i}^{2}-\alpha\tau^{i}\end{bmatrix}\succeq 0,$ (87a)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\alpha\geq 0.$ (87b)
Nevertheless, the above constraint would still be a nonconvex constraint, as ${\bf F}_{1,k}^{i}({\bf K}^{i})$, ${\bf F}_{2,k}^{i}({\bf K}^{i},\bar{{\bm{u}}}^{i})$, and ${\bf F}_{3,k}^{i}(\bar{{\bm{u}}}^{i})$ are nonlinear in ${\bf K}^{i}$, $\bar{{\bm{u}}}^{i}$. To address this issue, we first linearize the constraint (84) with respect to ${\bf K}^{i},\bar{{\bm{u}}}^{i}$ around the nominal ${\bf K}^{i,l},\bar{{\bm{u}}}^{i,l}$, to obtain the following affine robust constraint
@@
B-B Proposition 2 Proof
Let us initially rewrite the constraint ( 33 ) as follows